OpenStudy (senia01):

A line has a slope of and a y-intercept of −5. Which equations represent the line? Choose all answers that are correct. A. B. C. D. I already know B is right but I need more than one answer

OpenStudy (senia01):

A. -2x+3y=-15

OpenStudy (senia01):

B. y=2/3x-5

OpenStudy (senia01):

C. .6y=4x-30

OpenStudy (senia01):

D. (y+3)=2/3(x-3)

OpenStudy (senia01):

the slope is 2/3

OpenStudy (texaschic101):

slope (m) of 2/3 and y int of -5... -2x + 3y = -15 3y = 2x - 15 y = 2/3x - 5.....here we have slope of 2/3 and y int of 5...so not this one B is correct 6y = 4x - 30 y = 4/6x - 30/6 y = 2/3x - 5...here we have slope of 2/3 and y int of -5...so not this one y + 3 = 2/3(x - 3) y + 3 = 2/3x - 1/2 y = 2/3x - 1/2 - 3 y = 2/3x - 1/2 - 6/2 y = 2/3x - 7/2....here we have slope of 2/3 and y int of -7/2...not this one I believe there is only 1 answer and it is B

OpenStudy (texaschic101):

oops...that first one, I meant to say...slope = 2/3 and y int = -5...not this one

OpenStudy (texaschic101):

wait...I read this wrong...hold on

OpenStudy (senia01):

Thank you so much I've been stuck for 20 minutes.

OpenStudy (texaschic101):

It is A, B, and C

OpenStudy (senia01):

Thank you

OpenStudy (texaschic101):

sure thing...sorry about my mistake